# Scalarisation-based risk concepts for robust multi-objective optimisation

### v1

@@ 5 Robust performance metrics
Performance metrics are often used in multi-objective optimisation in order to quantify the quality of a Pareto front approximation. As showcased in a recent survey by Audet et al. (2021), there exists many different types of performance metrics in the multi-objective literature. All of which try to measure different facets of the Pareto front approximation. As motivated by Tu et al. (2024a), the most natural class of performance metrics, under the scalarisation perspective (Section 2.2) of multi-objective optimisation, are the family of R2 utilities (Hansen and Jaszkiewicz, 1998). Notably, this family of performance metrics exhibit many favourable properties and contains many of the most widely-used performance metrics as special cases (Tu et al., 2024a). In Section 5.1, we will first recall the definition of an R2 utility before generalising this concept to the robust setting via the RTS and STR methodology. Consequently, in the sections that follow, we give some concrete examples on how one can construct these robust adaptations in practice. Explicitly, we propose some robust generalisations to the well-known inverted generational distance (Section 5.2) and the hypervolume indicator (Section 5.3).
@@ 5.1 Robust R2 utilities
for any set of inputs $X\subseteq\mathbb{X}$. Note that in contrast to the original R2 utilities (22), these robust R2 utilities assess the quality of a set of inputs instead of a set of outputs. Formally, these robust performance metrics are defined with respect to an objective function and a risk functional. This construction makes sense from an intuitive perspective because the output of any input is subject to uncertainty and therefore one has to marginalise over this uncertainty in some way in order to assess the overall robust performance.
@@ 5.3 Robust hypervolume indicator
Figure 6: An illustration of the standard and robust IGD (Section 5.2) and hypervolume (Section 5.3) indicators on a two-dimensional example. In the plots, we illustrate the output associated with three inputs which are assumed to be either fixed for the standard examples or uniformly distributed on some ellipsoid for the robust examples. For the robust utilities, we used the expectation functionals; for the IGD indicator, we use three utopia points (stars); and for the HV indicator, we set the reference point to the origin (square).
@@ 6.1 Cake problem
In practice, given a data set such as this, one wishes to solve the inverse problem. That is, find me a cake whose properties achieves a certain set of desirable values. As this data set only contains a finite number of experiments, we follow standard procedure and fit a probabilistic model on it. Intuitively, this model is used as a surrogate in order to get a prediction on the properties of a cake at any feasible set of ingredients. For convenience, in this example, we fit a Gaussian process model, which is a popular model that is often used in industrial applications—see the book by Rasmussen and Williams (2006) for a review. More precisely, we let $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, denote the corresponding posterior Gaussian process model that was fitted on the data points. That is, every instance of the uncertain parameter $\boldsymbol{\xi}\in\Xi$ corresponds to a different sample path of the Gaussian process model:
@@
for any finite set of inputs $X=(\mathbf{x}_{1},\dots,\mathbf{x}_{N})\in\mathbb{X}^{N}$, where $\boldsymbol{\mu}_{T}:\mathbb{X}\rightarrow\mathbb{R}^{M}$ and $\boldsymbol{\Sigma}_{T}:\mathbb{X}\times\mathbb{X}\rightarrow\mathbb{R}^{M\times M}$ denotes the corresponding mean and covariance function of the posterior Gaussian process model. In Figure 8, we plot the posterior mean of this model over a sensible discretisation of the input space. We also highlight the corresponding Pareto optimal vectors and Pareto front surface.
@@
$\mathbf{x}^{*}_{\text{RTS}}$ $=(0.0323,0.2581,0.2903,0.0968,0.1935,0.1290)\in\mathbb{X},$
$\mathbf{x}^{*}_{\text{STR}}$ $=(0.0323,0.2258,0.2258,0.1613,0.1935,0.1613)\in\mathbb{X}.$
@@ 6.2 Robust rocket injector problem
To emulate a real-world scenario, we treat the objective function $f$ as a black-box function that can only be evaluated sparingly at a few carefully selected design locations $T=65$. Precisely, at any time $t>0$, the decision maker is given the opportunity to see a noisy function evaluation at any chosen design $\mathbf{x}_{t}\in\mathbb{X}$, that is $\mathbf{y}_{t}=f(\mathbf{x}_{t},\boldsymbol{\xi}_{t})+\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{M}$, where $\boldsymbol{\xi}_{t}\sim p(\boldsymbol{\xi})$ is the uncertain parameter that was independently sampled by the manufacturer and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{M},\textsc{Diag}(\boldsymbol{\sigma}^{2}))$ is the observation noise. In our example, we consider a zero-mean Gaussian observation noise whose standard deviation $\boldsymbol{\sigma}\approx(0.01,0.01,0.01)\in\mathbb{R}_{\geq 0}^{M}$ is set to $1\%$ of the estimated scaled objective range. To initialise this problem, we randomly sampled $T_{0}=5$ initial points from the controllable input space. In Figure 9, we illustrate these initial points in the objective space and plot the corresponding CVaR front under the STR approach. Notably, to compute this CVaR front, we require knowledge of the true objective function at each uncertain parameter which is typically not available in an actual real-world experimental set-up.
@@
In Figure 10, we present the results of our numerical comparison in which we compared our proposed method versus the existing baselines over $100$ independent runs on a discretised version of the robust rocket injector problem. Namely we compare our method against non-robust approaches such as: the random search (Sobol), NSGA2 (Deb et al., 2002), NParEGO (Knowles, 2006; Daulton et al., 2021), NEHVI (Emmerich et al., 2006; Daulton et al., 2021), PFES (Suzuki et al., 2020) and the JES-LB (Tu et al., 2022) algorithms. We also compare our approach against the MARS algorithm (Daulton et al., 2022), which is a robust algorithm that takes an STR approach using the Chebyshev scalarisation function and the univariate value-at-risk risk functional. Unsurprisingly we see that the robust methods perform better than the non-robust methods. Moreover, our method performs the best overall. For some visual intuition, we have included Figure 11 where we plot the final CVaR front obtained by each algorithm under the same random seed. Visually we see that most of the variation between these algorithms transpire around the first and second objectives. The existing approaches regularly struggle to find the inputs which robustly maximises the CVaR front over these two objectives. In contrast, our robust approach is much more consistent and is often very quick in identifying the difficult inputs which robustly maximises the whole CVaR front.
@@ 7 Conclusion
In this work, we studied the robust multi-objective optimisation problem from a computational perspective. We identified that the majority of existing robust multi-objective optimisation algorithms, in one way or another, rely on two key operations: robustification and scalarisation. These two procedures are in general not commutative and therefore the order that they are applied in has an effect on the resulting solutions that are identified and the final decisions that are made. As part of our work, we analysed the effect of this ordering from both a philosophical and a computational perspective. In particular, after setting up the problem in Sections 2 and 3, we presented a useful discussion on the different scenarios when one should prefer one ordering over another (Section 3.4.1) and when these approaches are equivalent (Section 3.4.2). We then moved on to the next sections where we described how these two different orderings can be used in order to extend classical concepts and tools from standard multi-objective optimisation to the robust setting. Specifically, we defined the general concept of a robust Pareto front (Section 4) and a robust performance metric (Section 5). We then demonstrated the utility of these new ideas via some insightful numerical experiments in Section 6.
@@ A.5 Pareto front conditions
(Tu et al., 2024b, Proposition 3.1) Consider a polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$, then this set is a Pareto front surface, $A=\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]$, if and only if the following two conditions holds:
@@
Suppose that the Pareto front surface of interest is non-empty. We can then invoke the polar parametrisation result (Theorem 4.1) on the Pareto front surfaces $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\cup_{\mathbf{x}\in X}\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]$ and $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\cup_{\mathbf{x}\in X}\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]]$. We see very quickly that both of these expressions reduce down to the RTS front (20) and therefore the latter set is indeed a Pareto front surface in this setting.
@@
This equivalence follows directly from the polar parametrised form of the risk statistic and the definition of the length scalarisation function. In essence it says that the point within the risk statistic which achieves the largest projected length along the line $\{\boldsymbol{\eta}+t\boldsymbol{\lambda}\in\mathbb{R}^{M}:t\in\mathbb{R}\}$ is indeed the vector $\boldsymbol{\eta}+\rho[s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(f(\mathbf{x},\cdot))]\boldsymbol{\lambda}\in\mathbb{R}^{M}$.
@@
Suppose that the Pareto front surface of interest is non-empty. We can then invoke the polar parametrisation result (Theorem 4.1) on this Pareto front surface. By using the equivalence in (31), we see very quickly that this polar representation reduces down to the STR front (21) and therefore the latter set is indeed a valid Pareto front surface in this setting.
@@ A.10 Hypervolume of a Pareto front surface
We now repeat a useful result in Lemma A.1, which states that the hypervolume of any Pareto front surface can be written as an integral over the set of positive unit vectors. This result follows simply from a change of coordinates. That is, we rewrite the volume integral using polar coordinates. A similar derivation was presented in the work by Deng and Zhang (2019) and Zhang and Golovin (2020) for the standard hypervolume indicator (27).
@@
A.11 Proof of Proposition 5.1
This result follows directly from the definition of the RTS front (20), together with Proposition 4.2, Lemma A.1 and the monotonicity of the hypervolume transformation.
@@
A.12 Proof of Proposition 5.2
This result follows directly from the definition of the STR front (21), together with Proposition 4.3, Lemma A.1 and the monotonicity of the hypervolume transformation.
$\blacksquare$

### v2

@@ 5 Robust performance metrics
Performance metrics are often used in multi-objective optimisation in order to quantify the quality of a Pareto front approximation. As showcased in a recent survey by Audet et al. (2021), there exists many different types of performance metrics in the multi-objective literature. All of which try to measure different facets of the Pareto front approximation. As established by Tu et al. (2024a), the most natural class of performance metrics, under the scalarisation perspective, are the family of R2 utilities originally proposed by Hansen and Jaszkiewicz (1998). Notably, this family of performance metrics exhibit many favourable properties and contains many of the most widely-used performance metrics as special cases (Tu et al., 2024a). We will now generalise this family of utilities to the robust setting via the RTS and STR methodology. In particular, we will illustrate some robust generalisations to the well-known inverted generational distance (Section 5.2) and the hypervolume indicator (Section 5.3).
@@ 5.1 Robust R2 utilities
for any set of inputs $X\subseteq\mathbb{X}$. Note that in contrast to the original R2 utilities in (22), these robust R2 utilities assess the quality of a set of inputs instead of a set of outputs. Formally, these robust performance metrics are defined with respect to an objective function and a risk functional. This construction makes sense from an intuitive perspective because the output of any input is subject to uncertainty and therefore one has to account for the uncertainty in some way in order to assess the overall robust performance.
@@ 5.3 Robust hypervolume indicator
Figure 6: An illustration of the standard and robust IGD (Section 5.2) and hypervolume (Section 5.3) indicators on a two-dimensional example. In the plots, we illustrate the output associated with three inputs which are assumed to be either fixed for the standard (no uncertainty) examples, or uniformly distributed on some ellipsoid for the robust examples. For the robust utilities, we used the expectation functionals; for the IGD indicator, we use three utopia points (stars); and for the HV indicator, we set the reference point to the origin (square).
@@ 6.1 Cake problem
In practice, given a data set such as this, one wishes to solve the inverse problem. That is, to find a cake whose properties achieves a certain set of desirable values. As this data set only contains a finite number of experiments, we follow standard procedure and fit a probabilistic model on it. Intuitively, this model is used as a surrogate in order to get a prediction on the properties of a cake at any feasible set of ingredients. For convenience, in this example, we fit a Gaussian process model, which is a popular model that is often used in industrial applications—see for instance Rasmussen and Williams (2006) for a review. More precisely, we let $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, denote the corresponding posterior Gaussian process model that was fitted on the data points. That is, every instance of the uncertain parameter $\boldsymbol{\xi}\in\Xi$ corresponds to a different sample path of the Gaussian process model:
@@
for any finite set of inputs $X=(\mathbf{x}_{1},\dots,\mathbf{x}_{N})\in\mathbb{X}^{N}$, where $\boldsymbol{\mu}_{T}:\mathbb{X}\rightarrow\mathbb{R}^{M}$ and $\boldsymbol{\Sigma}_{T}:\mathbb{X}\times\mathbb{X}\rightarrow\mathbb{R}^{M\times M}$ denotes the corresponding mean and covariance function of the posterior Gaussian process model. In Figure 8, we plot the posterior mean of this model over a discretisation of the input space with $|\hat{\mathbb{X}}|=44835$. We also highlight the corresponding Pareto optimal vectors and Pareto front surface.
@@
$\mathbf{x}^{*}_{\text{RTS}}$ $=(0.0323,0.2581,0.2903,0.0968,0.1935,0.1290),$
$\mathbf{x}^{*}_{\text{STR}}$ $=(0.0323,0.2258,0.2258,0.1613,0.1935,0.1613).$
@@ 6.2 Robust rocket injector problem
To emulate a real-world scenario, we treat the objective function $f$ as a black-box function that can only be evaluated sparingly at a few carefully selected design locations $T=65$. Precisely, at any time $t>0$, the decision maker is given the opportunity to see a noisy function evaluation at any chosen design $\mathbf{x}_{t}\in\mathbb{X}$, that is $\mathbf{y}_{t}=f(\mathbf{x}_{t},\boldsymbol{\xi}_{t})+\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{M}$, where $\boldsymbol{\xi}_{t}\sim p(\boldsymbol{\xi})$ is the uncertain parameter that was independently sampled by the manufacturer and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{M},\textsc{Diag}(\boldsymbol{\sigma}^{2}))$ is the observation noise. In our example, we consider a zero-mean Gaussian observation noise whose standard deviation $\boldsymbol{\sigma}=(0.01,0.01,0.01)$ is set to $1\%$ of the estimated scaled objective range. To initialise this problem, we randomly sampled $T_{0}=5$ initial points from the controllable input space. In Figure 9, we illustrate these initial points in the objective space and plot the corresponding CVaR front under the STR approach. Notably, to compute this CVaR front, we require knowledge of the true objective function at each uncertain parameter which is typically not available in an actual real-world experimental set-up.
@@
In Figure 10, we present the results of our numerical comparison in which we compared our proposed method versus the existing baselines over $100$ independent runs on a discretised version of the robust rocket injector problem. Namely we compare our method against non-robust approaches such as: the random search (Sobol), NSGA2 (Deb et al., 2002), NParEGO (Knowles, 2006; Daulton et al., 2021), NEHVI (Emmerich et al., 2006; Daulton et al., 2021), PFES (Suzuki et al., 2020) and the JES-LB (Tu et al., 2022) algorithms. We also compare our approach against the MARS algorithm (Daulton et al., 2022), which is a robust algorithm that takes an STR approach using the Chebyshev scalarisation function and the univariate value-at-risk risk functional. Unsurprisingly we see that the robust methods perform better than the non-robust methods. Moreover, our method performs the best overall under the chosen utility. For some visual intuition, we have included Figure 11 where we plot the final CVaR front obtained by each algorithm under the same random seed. Visually we see that most of the variation between these algorithms transpire around the first and second objectives. The existing approaches regularly struggle to find the inputs which robustly maximises the CVaR front over these two objectives. In contrast, our robust approach is much more consistent and is often very quick in identifying the difficult inputs which robustly maximises the whole CVaR front.
@@ 7 Conclusion
In this work, we studied the robust multi-objective optimisation problem from a computational perspective. We identified that the majority of existing robust multi-objective optimisation algorithms, in one way or another, rely on two key operations: robustification and scalarisation. These two procedures are in general not commutative and therefore the order that they are applied in has an effect on the resulting solutions that are identified and the final decisions that are made. As part of our work, we analysed the effect of this ordering from both a philosophical and a methodological perspective. In particular, after setting up the problem in Sections 2 and 3, we presented a useful discussion on the different scenarios when one should prefer one ordering over another (Section 3.3.2) and when these approaches are equivalent (Section 3.3.3). We then moved on to the next sections where we described how these two different orderings can be used in order to extend classical concepts and tools from standard multi-objective optimisation to the robust setting. Specifically, we defined the general concept of a robust Pareto front (Section 4) and a robust performance metric (Section 5). We then demonstrated the utility of these new ideas via some insightful numerical experiments in Section 6.
@@ A.5 Pareto front conditions
(Tu et al., 2024b, Proposition 3.1) Consider a polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$, then this set is a Pareto front surface, that is $A=\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]$, if and only if the following two conditions holds:
@@
Suppose that the Pareto front surface of interest is non-empty. We can then invoke Theorem 4.1 on the Pareto front surfaces $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\cup_{\mathbf{x}\in X}\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]$ and $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\cup_{\mathbf{x}\in X}\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]]$. We see very quickly that both of these expressions reduce down to the RTS front (20) and therefore the latter set is indeed a Pareto front surface in this setting.
@@
This equivalence follows directly from the polar parameterised form of the risk statistic and the definition of the length scalarisation function. In essence it says that the point within the risk statistic which achieves the largest projected length along the line $\{\boldsymbol{\eta}+t\boldsymbol{\lambda}\in\mathbb{R}^{M}:t\in\mathbb{R}\}$ is indeed the vector $\boldsymbol{\eta}+\rho[s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(f(\mathbf{x},\cdot))]\boldsymbol{\lambda}\in\mathbb{R}^{M}$.
@@
Suppose that the Pareto front surface of interest is non-empty. We can then invoke Theorem 4.1 on this Pareto front surface. By using the equivalence in (31), we see very quickly that this polar representation reduces down to the STR front (21) and therefore the latter set is indeed a valid Pareto front surface in this setting.
@@ A.10 Hypervolume of a Pareto front surface
We now recall a useful result in Lemma A.1, which states that the hypervolume of any Pareto front surface can be written as an integral over the set of positive unit vectors. A similar result was presented in the work by Deng and Zhang (2019) and Zhang and Golovin (2020) for the standard hypervolume indicator (27).
@@
A.11 Proof of Proposition 5.1 and Proposition 5.2
The proof of Proposition 5.1 follows directly from the definition of the RTS front (20), together with Proposition 4.2, Lemma A.1 and the monotonicity of the hypervolume transformation. The proof of Proposition 5.2 follows in a similar way.